Sol Summary – April 19th

Crew 280 Sol Summary Report 19-04-2023

Sol: 3

Summary title: Indoor Adventures for a Windy Day on Mars

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary: The Hypatia I crew woke up with high spirits and a renewed sense of adventure. As the crew began their day, some members attended a yoga class taught by Ariadna Farrés (Health & Safety Officer), while others had the chance to take their first shower in days. The refreshing feeling of clean skin was a welcome change from the usual routine of sponge baths and hygiene wipes!

Once everyone was ready for the day, the crew gathered on the upper deck to have breakfast and attend a group meeting. Originally, the Hypatia I crew had planned to perform an EVA to Compass Rock with three main goals in mind: (i) collect sand samples for Laia Ribas’ (GreenHab Officer) CatMart project, (ii) evaluate the feasibility of different navigation strategies developed by Ariadna Farrés (Scientist & HSO); and (iii) test our own Midland radios for long-distance communication. However, as they checked the forecast, they realized that high winds were expected throughout the day. Therefore, the crew decided to be conservative and cancel their EVA. It was a tough call, as they were excited to work on their experiments outside the habitat. But safety always comes first on Mars, and they did not want to risk the well-being of anyone on the mission.

As the afternoon progressed, the crew took advantage of their time and focused on their experiments inside the habitat. At some point, the heater stopped working, causing the Hab to become quite cold. Cesca Cufí (Crew Engineer) was able to troubleshoot this problem with the assistance of Mission Support, but it is now failing again. We hope to be able to fix it, as temperatures will continue to drop during the night.

Despite the unexpected challenge, the crew remained positive, looking forward to a relaxing evening ahead. They will have homemade bread for dinner and are now excited to wind down with a delicious meal and a movie, ready for whatever challenges that tomorrow might bring.

Look ahead plan: We are planning on conducting two EVAs and continuing our experimental work.

Anomalies in work: Non-functional robotic telescope and Spirit needs a new battery wire.

Weather: Sunny, but high winds. Low 37F/High 53F.

Crew physical status: Nominal

EVA: Not applicable today.

Reports to be filed: Sol Summary, EVA request, HSO Report, Operations Report, GreenHab Report, Journalist Report, Pictures of the Day.

Support Requested: Replacement of stove knobs, if possible.

Copyright © The Mars Society. All rights reserved. | Main Site